# What is the effective size of a parity-time symmetric dimer?

### 02 Jan 2024 v1

@@ Abstract
Parity-time (PT) symmetric dimers were introduced to highlight the unusual properties of non-Hermitian systems that are invariant after a combined parity and time reversal operation. They are also the building blocks of a variety of symmetry and topologically protected structures, especially on integrated photonic platforms. As the name suggests, it consists of two coupled oscillators, which can be optical, mechanical, electronic, etc. in nature. In this article, we show that its effective size, defined by the number of lattice sites inversely proportional to the lattice momentum, is surprisingly three instead of two from the perspective of energy quantization. More specifically, we show analytically that the complex energy levels of a one-dimensional concatenated chain with $N$ PT-dimers are determined by a system size of $1+2N$, which reduces to three in the case of a single PT-dimer. We note that while energy quantization conditions have been established in various non-Hermitian systems, exact and explicitly quantized complex energies as reported here are still scarce. In connection, we also discuss the other symmetries of a PT-dimer and concatenated PT-dimer chain, including non-Hermitian particle-hole symmetry and chiral symmetry.
@@ I Introduction
In this article, we focus on a different and seemingly apparent question: What is the effective size of a PT-symmetric dimer, defined by the number of lattice sites inversely proportional to the lattice momentum? This question is important in understanding the discretized energy levels of a finite-sized system, whether they are Hermitian or non-Hermitian. Given that a PT-dimer consists of two coupled oscillators, one may immediately jump to the conclusion that the answer is two. However, as we will show below, its appropriate size is three instead of two from the perspective of energy quantization. This statement, of course, requires clarification, as one does not observe a factor of three (or even two) in the two energy eigenvalues of a PT-dimer. However, once we consider a concatenation of $N$ PT-dimers with the same coupling, a crucial factor given by $1+2N$ emerges that represents the size of the chain, which becomes three in the case of a single PT-dimer.
We note that while energy quantization conditions for non-Hermitian systems have been established in various scenarios, including high-precision approximations [2], an imaginary coordinate shift of exactly solvable Hermitian potentials [20], and the Milne quantization and supersymmetry algebra [21], exact and explicitly quantized complex energies as reported here are still scarce. In fact, one may have wondered whether such quantized complex energies, determined by an integer quantum number, exist at all in a simple non-Hermitian system, and our finding provides an affirmative answer to this question. We also note that although most PT-dimers and chains demonstrated so far experimentally are classical (see, e.g., Fig. 1 and Ref. [7]), the energy quantization procedure is the same as in non-interacting quantum systems due to their shared wave nature. For example, the evanescant coupling between two optical elements plays the same role as the tunneling of electrons between two atomic orbitals, and hence the respective energy bands they induce can be treated on the same footing. In connection, we also discuss the other symmetries of a PT-dimer and the aforementioned PT-dimer chain, including non-Hermitian particle-hole symmetry [22, 23, 24, 25, 26] and chiral symmetry [27].
Figure 1: Schematics and representative integrated photonic realizations of a PT-dimer (a) and a concatenation of PT-dimers. The experimental figures are adapted from Refs. [3, 4, 5, 6, 19, 18, 17, 16] with permissions. We take all couplings in (b) to be the same in the main text.
@@ III Energy quantization
where $\theta\equiv\cos^{-1}\alpha$ is real (imaginary) if $|\alpha|<1$ ($|\alpha|>1$). By substituting equation (14) with $m=1,2$ (i.e., for $\Psi_{1},\Psi_{3}$) into equation (10), we find
@@
In the last step, we have used $\alpha=\cos\theta$ and the identity
@@
similar to equation (9). Using $n=2N-1$ (and $m=N$) in Eq. (17) and we find $\Psi_{2N-1}=e^{iN\theta}-e^{-i(N-1)\theta}$, which gives
@@
As we have mentioned in the caption of Fig. 2, the spacing between two sites in a PT-dimer is taken to be 1 and dimensionless, and from the energy quantization condition above, we then find
@@
to be the effective size of the concatenated PT-dimer chain. For a single PT-dimer, $N=1$ and hence its effective size is three instead of two in this regard. To verify that equation (27) still holds in this case, we observe that it reduces correctly to Eq. (2) as anticipated, i.e.,
@@ V Multiple symmetries
When introducing $\alpha$ in Eq. (13), we have mentioned that $\alpha$ is real using the property that $\varepsilon_{\mu}$’s must be real or imaginary as determined by the symmetries of the system. $\varepsilon_{\mu}$’s are real in the PT-symmetric phase, but PT symmetry does not warrant them to be imaginary in the PT-broken phase. This observation has not been appreciated in non-Hermitian photonics with quantum-inspired symmetries, and there is indeed another symmetry that pins the complex energy levels on the imaginary axis in the PT-broken phase.
@@
As we have seen in Sec. III, these symmetries do not actually play a role in our derivation of the energy quantization condition. Nevertheless, they do make our result easier to interpret as both $\alpha$ and $\theta$ are real. Interestingly, even if we lift PT symmetry (and consequently, also non-Hermitian chiral symmetry) explicitly by removing the leftmost site in Fig. 1b, the energy levels of our concatenated chain are still either real or imaginary, as we demonstrate below.
@@
In other words, they are still either real or imaginary as we have mentioned (see Figs. 3(d-f)). Furthermore, they are still paired with just a sign difference except for $\varepsilon_{2N-1}$.
Using the momentum $k$ in place of $\theta$ as we did before, we then write the energy quantization condition as
@@ VI Discussions
In summary, we have discussed the effective size of concatenated PT-dimer chains and the symmetry properties of their spectra, both with and without a boundary defect by removing a boundary site. In the limit that the number of PT-dimers goes to one, we find that the effective size of this single PT-dimer is three instead of two, despite the absence of apparent size-related factors in the energy levels. Similarly, when there is one and half PT-dimers, we find that its effective size is four instead of three from Eq. (44).

### 14 Feb 2024 v2

# Effective size of a parity-time symmetric dimer

@@ Abstract
Parity-time (PT) symmetric dimers were introduced to highlight the unusual properties of non-Hermitian systems that are invariant after a combined parity and time reversal operation. They are also the building blocks of a variety of symmetry and topologically protected structures, especially on integrated photonic platforms. As the name suggests, it consists of two coupled oscillators, which can be optical, mechanical, electronic, and so on in nature. In this article, we show that its effective size, defined by the number of lattice sites inversely proportional to the lattice momentum, is surprisingly three instead of two from the perspective of energy quantization. More specifically, we show analytically that the complex energy levels of a one-dimensional concatenated chain with $N$ PT-dimers are determined by a system size of $1+2N$, which reduces to three in the case of a single PT-dimer. We note that while energy quantization conditions were established in various non-Hermitian systems, exact and explicitly quantized complex energies as reported here are still scarce. In connection, we also discuss the other symmetries of a PT-dimer and concatenated PT-dimer chain, including non-Hermitian particle-hole symmetry and chiral symmetry.
@@ I Introduction
In this article, we focus on a different and seemingly apparent question: What is the effective size of a PT-symmetric dimer, defined by the number of lattice sites inversely proportional to the lattice momentum? This question is important in understanding the discretized energy levels of a finite-sized system, whether it is Hermitian or non-Hermitian. Given that a PT-dimer consists of two coupled oscillators, one may immediately jump to the conclusion that the answer is two. However, as we will show below, its appropriate size is three instead of two from the perspective of energy quantization. This statement, of course, requires clarification, as one does not observe a factor of three (or even two) in the two energy eigenvalues of a PT-dimer. However, once we consider a concatenation of $N$ PT-dimers with the same coupling, a crucial factor given by $1+2N$ emerges that represents the size of the chain, which becomes three in the case of a single PT-dimer.
We note that while energy quantization conditions for non-Hermitian systems have been established in various scenarios, including high-precision approximations [2], an imaginary coordinate shift of exactly solvable Hermitian potentials [20], and the Milne quantization and supersymmetry algebra [21], exact and explicitly quantized complex energies as reported here are still scarce. In fact, one may have wondered whether such quantized complex energies, determined by an integer quantum number, exist at all in a simple non-Hermitian system, and our finding provides an affirmative answer to this question. We also note that although most PT-dimers and chains demonstrated so far experimentally are classical (see, e.g., Fig. 1 and Ref. [7]), the energy quantization procedure is the same as in non-interacting quantum systems due to their shared wave nature. For example, the evanescent coupling between two optical elements plays the same role as the tunneling of electrons between two atomic orbitals, and hence the respective energy bands they induce can be treated on the same footing. In connection, we also discuss the other symmetries of a PT-dimer and the aforementioned PT-dimer chain, including non-Hermitian particle-hole symmetry [22, 23, 24, 25, 26] and chiral symmetry [27].
Figure 1: Schematics and representative integrated photonic realizations of (a) a PT-dimer and (b) a concatenation of PT-dimers. The experimental figures are adapted from Refs. [3, 4, 5, 6, 16, 17, 18, 19] with permissions. We take all couplings in (b) to be the same in the main text.
@@ III Energy quantization
where $\theta\equiv\cos^{-1}\alpha$ is real (imaginary) if $|\alpha|<1$ ($|\alpha|>1$). By substituting Eq. (14) with $m=1,2$ (i.e., for $\Psi_{1},\Psi_{3}$) into Eq. (10), we find
@@
In the last step above, we used $\alpha=\cos\theta$ and the identity
@@
similar to Eq. (9). Using $n=2N-1$ (and $m=N$) in Eq. (17) and we find $\Psi_{2N-1}=e^{iN\theta}-e^{-i(N-1)\theta}$, which gives
@@
As we mentioned before, the spacing between two sites in a PT-dimer is taken to be 1 and dimensionless, and from the energy quantization condition above, we then find
@@
to be the effective size of the concatenated PT-dimer chain. For a single PT-dimer, $N=1$ and hence its effective size is three instead of two in this regard. To verify that Eq. (27) still holds in this case, we observe that it reduces correctly to Eq. (2) as anticipated, i.e.,
@@ V Multiple symmetries
When introducing $\alpha$ in Eq. (13), we mentioned that $\alpha$ is real using the property that $\varepsilon_{\mu}$’s must be real or imaginary as determined by the symmetries of the system. $\varepsilon_{\mu}$’s are real in the PT-symmetric phase, but PT symmetry does not warrant them to be imaginary in the PT-broken phase. This observation has not been appreciated in non-Hermitian photonics with quantum-inspired symmetries, and there is indeed another symmetry that pins the complex energy levels on the imaginary axis in the PT-broken phase.
@@
As we saw in Sec. III, these symmetries do not actually play a role in our derivation of the energy quantization condition. Nevertheless, they do make our result easier to interpret as both $\alpha$ and $\theta$ are real. Interestingly, even if we lift PT symmetry (and consequently, also non-Hermitian chiral symmetry) explicitly by removing the leftmost site in Fig. 1b, the energy levels of our concatenated chain are still either real or imaginary, as we demonstrate below.
@@
In other words, they are still either real or imaginary as we mentioned (see Figs. 3(d-f)). Furthermore, they are still paired with just a sign difference except for $\varepsilon_{2N-1}$.
We mention in passing that this difference from the PT-symmetric structure leads to a dramatic change of the eigenvalues of the scattering matrix [33], which connects the system to incoming and outgoing channels. More specifically, PT symmetry warrants that when a pole of the scattering matrix reaches the real frequency axis, it coincides with a zero of the same scattering matrix [34]. This property leads to a CPA-laser [35] where CPA stands for coherent perfect absorption [36, 37]. Interestingly, removing one loss layer from one end of a PT-symmetric optical heterostructure (corresponding to our removal of the leftmost site when $\gamma$ is negative) does not affect the operation of a CPA-laser [38], but the linewidth displayed by the eigenvalues of the scattering matrix can change by orders of magnitude [33].
To find the effective size of the non-PT chain here with $12N-$ 1 sites, we use the momentum $k$ in place of $\theta$ as before, and we write the energy quantization condition as
@@ VI Discussions
In summary, we discussed the effective size of concatenated PT-dimer chains and the symmetry properties of their spectra, both with and without a boundary defect by removing a boundary site. In the limit that the number of PT-dimers goes to one, we find that the effective size of this single PT-dimer is three instead of two, despite the absence of apparent size-related factors in the energy levels. Similarly, when there is one and half PT-dimers, we find that its effective size is four instead of three from Eq. (44).
@@
Acknowledgements.
This work is supported by the National Science Foundation (NSF) under grant No. PHY-1847240.